OpenStudy (anonymous):

(d/dx)e^(lny) = dy/dx,... why does this not equal e^(lny)(dy/dx) by chain rule?

OpenStudy (anonymous):

why does it not equal (y)(dy/dx)

OpenStudy (anonymous):

set u = ln y du = dy/y y = e^ln y = e^ln u d(e^lny)/dx = (d(e^u)/du)(du/dx) = y. d(e^u) / dy = y(dy/y)/dx = dy/dx note: d(e^u) = d(e^lny) = dy

OpenStudy (anonymous):

y(dy/dx)=dy/dx??

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%28dy%2Fdx%29

hartnn (hartnn):

\(\huge x^{\log_x a}=a\)

hartnn (hartnn):

so, \(\huge e^{\ln y}=y\)

hartnn (hartnn):

d/dx (e^{ln y}) = d/dx (y) = dy/dx

OpenStudy (anonymous):

yeah, but still the chain rule should apply in the same way, that it assumes the dirivative of x^2 = 2x*(d/dx(x))= 2x(1)= 2x

OpenStudy (anonymous):

it should be that (d/dx)e^(lny)=e^(lny)*dy/dx = y(dy/dx)

OpenStudy (anonymous):

(*assuming you are using the chain rule definition, which i understand no reason to say that you couldn't)

hartnn (hartnn):

applying chain rule \(\large d/dx [e^{\ln y}]=e^{\ln y}d/dx[\ln y] = e^{\ln y}(1/y)d/dx[y]=e^{\ln y}(1/y)dy/dx\)

hartnn (hartnn):

(d/dx)e^(lny)=e^(lny)*dy/dx <------NO (d/dx)e^(lny)=e^(lny)*d[ln y]/dx

hartnn (hartnn):

\(\large d/dx [e^{\ln y}]=e^{\ln y}d/dx[\ln y] \\ = e^{\ln y}(1/y)d/dx[y]=e^{\ln y}(1/y)dy/dx \\ =(y)(1/y)dy/dx = dy/dx\)

hartnn (hartnn):

you get same answer using chain rule also.

OpenStudy (anonymous):

ok the foundations of mathematics still hold, for now ... XD, i see where i have gone wrong thank you for your explination! :)

hartnn (hartnn):

welcome ^_^